### v1

# Maximum Quantum Work at Criticality: Stirling Engines and Fibonacci-Lucas Degeneracies

Abstract
Many-body effects and quantum criticality play a central role in determining the performance of quantum thermal machines. Although operating near a quantum critical point (QCP) is known to enhance engine performance, the precise thermodynamic conditions required to attain the Carnot efficiency limit remain unsettled. Here, we derive the exact conditions for a quantum Stirling engine to achieve Carnot efficiency when a QCP drives its working medium. In the low-temperature regime, where only the ground-state manifold is populated, the net work output is given by $W=k_{B}\delta\ln(g_{\text{crit}}/g_{0})$ with $\delta=T_{H}-T_{L}$, which directly yields the Carnot efficiency $\eta_{C}=1-T_{L}/T_{H}$, independent of microscopic details. Notably, whereas ideal Stirling cycles attain Carnot efficiency only with a perfect regenerator, here no regenerator is required because, at low temperatures, the thermal population remains confined to the degenerate ground state; this represents a clear quantum advantage over engines with classical working substances. We validate this universal result by recovering known behaviors in various quantum systems, including spin chains with Dzyaloshinskii–Moriya interactions and magnetic anisotropies. Applying the framework to the one-dimensional antiferromagnetic Ising model, we predict non-extensive scaling of the work output governed by Fibonacci and Lucas numbers for open chains and closed rings, respectively, which converges to classical extensivity in the thermodynamic limit. This analysis establishes a general and robust foundation for designing quantum thermal machines that reach the Carnot bound while delivering finite work.
I Introduction
Quantum thermodynamics investigates how quantum principles govern energy exchange and conversion at microscopic scales [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. A central objective is to develop and analyze quantum thermal machines such as engines, refrigerators, and heat pumps operating in the quantum regime that may exploit coherence, entanglement, many-body effects, flat bands, and others to surpass classical performance benchmarks [14, 15, 16, 1, 17].
Among quantum cycles, the Stirling cycle has attracted sustained interest due to its conceptual simplicity and experimental accessibility to obtain finite work with highly efficiency [18, 19, 20, 21]. It consists of two isothermal and two isoparametric (or isochoric) processes, providing a versatile setting to probe performance limits. A recurring observation is that tuning the working medium near a quantum critical point (QCP) can markedly enhance performance [22, 23, 24, 25, 26, 27]. A QCP triggered at zero temperature by varying an external control parameter, such as field or pressure, features qualitative ground-state changes, divergent correlation lengths, and universal scaling [28]. The associated large density of states and enhanced fluctuations are widely regarded as a resource for thermodynamic tasks.
Extensive studies across diverse platforms, including Heisenberg spin chains [29, 30, 31, 32], anisotropic qubit systems [33, 34], and models with Dzyaloshinskii–Moriya interactions [35, 36, 37, 38], have reported enhanced work output and efficiency when thermodynamic cycles operate across a QCP. Nevertheless, a general and exact analytical characterization of the necessary and sufficient conditions to attain Carnot efficiency in critical quantum Stirling engines has remained open, with many results relying on numerics or model-specific features, leaving questions about universality and the role of ground-state degeneracies.
Here we present a general analytical framework for Stirling engines operating at a QCP. We show that, in the low-temperature and large-gap regime, a cycle executed between a critical point with degeneracy $g_{\text{crit}}$ and a high-field point with degeneracy $g_{0}$ achieves precisely Carnot efficiency, with net work given by a simple universal expression that depends only on the reservoir temperatures and the degeneracy ratio. We validate this result by quantitative agreement with previously reported systems ranging from two-qubit setups to extended spin chains with nontrivial interactions. We further apply the framework to the one-dimensional antiferromagnetic Ising model [39, 40, 41], unveiling a connection to number theory: the work scales with the logarithm of Fibonacci numbers for open chains and Lucas numbers for rings, revealing a non-extensive quantum effect that disappears only in the macroscopic limit. Notably, this critical-cycle route to Carnot efficiency does not require a classical regenerator.
In summary, our framework consolidates diverse findings into a unified theory and establishes experimentally testable conditions for quantum thermal machines to attain the fundamental efficiency limit with finite work output.
II Model
[…]
Figure 1: Schematic diagram of energy $E$ versus the tuning parameter $\lambda$. At the QCP located at $\lambda_{\mathrm{crit}}$, the $g_{\mathrm{crit}}$-fold degenerate ground state energy levels separate. The new ground state $\epsilon_{0}$ with degeneracy $g_{0}$ and the first excited state $\epsilon_{1}$ with degeneracy $g_{1}$ are separated by an energy gap $\Delta$ at high tuning parameter values $\lambda_{H}$. The thermal populations of higher excited states (with energies $\epsilon_{n>1}$) are negligible near absolute zero; only their degeneracies at the QCP influence the thermodynamic properties of the system.
The working medium consists of any physical system whose energy spectrum can be tuned by an external parameter $\lambda$, and which exhibits a QCP at $\lambda=\lambda_{\mathrm{crit}}$ when the temperature approaches absolute zero. At the QCP, the system undergoes a fundamental transformation characterized by symmetry breaking and qualitative changes in the ground state structure.
As illustrated in Fig. 1, the $g_{\mathrm{crit}}$-fold degenerate ground state at the QCP splits due to interactions with the tuning parameter. For $\lambda>\lambda_{\mathrm{crit}}$, the lowest-energy states separates into a ground state $\epsilon_{0}$ with degeneracy $g_{0}$ and an excited level $\epsilon_{1}$ characterized by degeneracy $g_{1}$. At sufficiently large values of the tuning parameter ($\lambda_{H}$), these two levels are separated by an energy gap $\Delta$.
Higher excited states $\epsilon_{n>1}$ play a negligible role in this regime. Near absolute zero temperature, the thermal population of these states becomes vanishingly small due to the Boltzmann suppression factor. Even at the QCP, the available thermal energy remains insufficient to significantly populate levels beyond the critic ground state. Only the degeneracies of these higher energy levels at $\lambda_{\mathrm{crit}}$ can influence the thermodynamic properties of the system through their contribution to the density of states at the critical point.
The thermodynamic properties of the system can be calculated and fully characterized through the canonical partition function. Taking into account that at absolute zero each energy level $E_{i}$ may be $g_{i}$-fold degenerate, the partition function is given by:
$Z_{\lambda}(T)=\sum_{i}g_{i}\exp\left(-\frac{E_{i}(\lambda)}{k_{B}T}\right)$ (1)
The thermal equilibrium populations follow directly from the partition function through the Boltzmann distribution:
$p_{i}(\lambda)\equiv\frac{g_{i}e^{-\beta E_{i}(\lambda)}}{Z_{\lambda}},\quad\text{where}\quad\beta=\frac{1}{k_{B}T}.$ (2)
From the partition function and the associated Boltzmann probabilities, the standard relations of canonical ensemble thermodynamics yield the Helmholtz free energy, entropy, and internal energy:
$F_{\lambda}(T)$ $=-k_{B}T\ln Z_{\lambda}(T),$ (3)
[…]
It is crucial to emphasize that in the zero-temperature limit for the two tuning parameters $\lambda_{\text{crit}}$ and $\lambda_{H}$ shown in Fig. 1, the internal energy reduces to the ground state energy, while the entropy, in analogy to the microcanonical ensemble formulation, is determined by counting the number of states sharing the same minimum energy, that is, the degeneracy of the ground state at each parameter value:
$U_{\lambda_{\text{crit}}}(T\rightarrow 0)=\epsilon_{\text{ground}}^{\text{(crit)}}$ $\implies$ $S_{\lambda_{\text{crit}}}(T\rightarrow 0)=k_{B}\ln\quantity(g_{\text{crit}})$ (6)
$U_{\lambda_{H}}(T\rightarrow 0)=\epsilon_{0}$ $\implies$ $S_{\lambda_{H}}(T\rightarrow 0)=k_{B}\ln\quantity(g_{0})$
III Stirling cycle and Criticallity
III.1 General Theory of the Quantum Stirling Cycle
The Stirling cycle is characterized by four stages: two isothermal processes and two isoparametric processes (where the tuning parameter $\lambda$ is held constant). Unlike the well-known Otto cycle, the net heat and work contributions in each segment are not directly evident from the trajectory of the cycle alone. Instead, they must be determined by analyzing the entropy differences at each stage point throughout the complete cycle.
The isoparametric stages are analogous to the isochoric processes in a classical ideal-gas Stirling cycle, where volume is fixed. In quantum and condensed matter systems, the fixed parameter $\lambda$ may represent various external controls such as magnetic field, electric field, exchange interactions, or mechanical stress, among others. During the isothermal processes, the system remains in thermal equilibrium with the hot and cold reservoirs at temperatures $T_{H}$ and $T_{L}$ ($T_{H}>T_{L}$), while the tuning parameter varies between $\lambda_{L}$ and $\lambda_{H}$ $(\lambda_{H}>\lambda_{L})$.
[…]
Figure 2: Schematic representation of the Stirling cycle in terms of entropy and the control parameter $\lambda$. Segments ABAB and CDCD correspond to the isothermal processes at $T=T_{H}$ and $T=T_{L}$, respectively. The entropy at point AA ($D$) is lower than at point $B$ ($C$), and the cycle proceeds in a counterclockwise direction.
During the isoparametric stages, the system undergoes temperature exchange at fixed parameter value. Consequently, no work is performed since the tuning parameter remains constant. The cycle representation in Fig. 2 will be analyzed stage-by-stage. It should be noted that any other intrinsic parameters of the system remain fixed throughout the cycle $\text{A}\rightarrow\text{B}\rightarrow\text{C}\rightarrow\text{D}\rightarrow\text{A}$; only the temperature $T$ and tuning parameter $\lambda$ are varied.
$\mathbf{A\rightarrow B}$.
Isothermal compression at $T=T_{H}$. The tuning parameter varies from $\lambda_{H}$ to $\lambda_{L}$ while the system maintains thermal equilibrium with the hot reservoir. The heat exchange during this process is given by the entropy change:
$Q_{AB}=T_{H}\quantity[S_{\lambda_{L}}(T_{H})-S_{\lambda_{H}}(T_{H})]$ (7)
$\mathbf{B\rightarrow C}$
Isoparametric cooling at constant $\lambda=\lambda_{L}$. Decoupling from the hot reservoir and cooling to $T_{L}$ decreases the internal energy without work performance:
$Q_{BC}=U_{\lambda_{L}}(T_{L})-U_{\lambda_{L}}(T_{H})$ (8)
$\mathbf{C\rightarrow D}$.
Isothermal expansion at $T=T_{L}$. The tuning parameter returns from $\lambda_{L}$ to $\lambda_{H}$ while the system equilibrates with the cold reservoir. The associated heat transfer is:
$Q_{CD}=T_{L}\quantity[S_{\lambda_{H}}(T_{L})-S_{\lambda_{L}}(T_{L})]$ (9)
$\mathbf{D\rightarrow A}$
Isoparametric heating at constant $\lambda=\lambda_{H}$. Heating from $T_{L}$ to $T_{H}$ increases the internal energy, again without work being performed:
$Q_{DA}=U_{\lambda_{H}}(T_{H})-U_{\lambda_{H}}(T_{L})$ (10)
From the first law of thermodynamics, the internal energy variation over a closed cycle is zero:
[…]
For this cycle, the total net work equals the sum of heat exchanges during each stage:
$W=Q_{AB}+Q_{BC}+Q_{CD}+Q_{DA}=Q_{H}+Q_{L}.$ (12)
We adopt the sign convention where positive (negative) heat $Q$ indicates that the system absorbs (releases) heat. For the cycle shown in Fig. 2, the net heat exchanges are $Q_{H}=Q_{AB}+Q_{DA}>0$ (absorbed from hot reservoir) and $Q_{L}=Q_{CD}+Q_{BC}<0$ (released to cold reservoir). Similarly, positive (negative) work $W$ indicates work done by (on) the system.
In the context of a heat engine performing useful work, the Stirling engine efficiency $\eta$ is defined as the ratio between the total net work and the heat absorbed from the hot reservoir:
$\eta=\frac{Q_{H}+Q_{L}}{Q_{H}}=\frac{W}{Q_{H}}=1-\left|\frac{Q_{CD}+Q_{BC}}{Q_{AB}+Q_{DA}}\right|.$ (13)
Depending on the system’s fundamental properties, the signs of work and heat may differ from those of a conventional heat engine. By reversing the cycle direction, a refrigerator can be obtained where previously there was an engine. Similarly, by applying substantial work to the system, it can extract heat from the cold reservoir and function as a heat pump. Furthermore, within the formalism of quantum thermodynamics, two additional exotic regimes exist: the accelerator and the heater. All these operational modes are summarized in the heat-work sign table in Table 1.
Table 1: Signs of heat and total work used to classify different operational regimes.
Operational regime $W$ $Q_{\mathrm{in}}$ $Q_{\mathrm{out}}$
Engine >0>0 >0>0 <0<0
Refrigerator <0<0 <0<0 >0>0
Heat pump <0<0 >0>0 >0>0
Heater <0<0 <0<0 <0<0
Accelerator <0<0 >0>0 <0<0
III.2 Quantum Stirling Engines at Quantum Criticality
While the general Stirling cycle shown in Fig. 2 works effectively for a system at quantum criticality as in Fig. 1, a specific diagram is necessary to illustrate the exact behavior. As illustrated in Fig. 3, the thermodynamic cycle is performed by sweeping the control parameter from the QCP $\lambda_{\text{crit}}$ (Fig. 1) to a high parameter value $\lambda_{H}$, with $\lambda_{L}=\lambda_{\text{crit}}$. Alternatively, by choosing to vary the high parameter $\lambda_{H}$ while keeping $\lambda_{L}$ fixed, the diagram becomes reflected horizontally, with point $A$ corresponding to the ground state at the higher temperature. Similarly, if the high field is selected as the control parameter, the cycle can be defined by assigning either the initial or the final point to coincide with the QCP.
[…]
Figure 3: Stirling cycle diagram for a system with a QCP, represented in terms of entropy and the control parameter $\lambda$. The critical value $\lambda_{\text{crit}}$ denotes the QCP. Segments ABAB and CDCD correspond to the isothermal processes at $T_{H}=T_{L}+\delta$ and $T_{L}$, respectively. The entropy satisfies $S_{B}=S_{C}=g_{\text{crit}}$, while $S_{D}=g_{0}$ and $S_{A}$ may include contributions from $g_{1}$. The heat exchange along BCBC vanishes ($Q_{BC}=0$), whereas $Q_{DA}$ is proportional to the temperature difference between reservoirs.
The thermal reservoirs are considered to be close in temperature, with $(T_{H}-T_{L})/T_{H}=\delta/T_{H}\ll 1$, so that $T_{H}$ may be written as $T_{H}=T_{L}+\delta$ to leading order. This condition ensures that when the system is at the QCP, the entropy is dominated by the ground-state degeneracy, with negligible population of excited states. Consequently, the internal energy at the critical point becomes temperature-independent and constant for both reservoir temperatures, since they are both small and closely spaced.
Additionally, the low-temperature regime guarantees that at high tuning parameter $\lambda_{H}$, predominantly the ground state is populated, while the first excited state remains scarcely occupied and higher levels are empty. The energy gap $\Delta$ between the ground and first excited states determines the characteristic energy scale, and the internal energy at both low and high temperatures is governed by this parameter. Later in this section, the case of a reduced energy gap will be considered, where thermal occupation of the first excited state becomes non-negligible even at low temperatures.
For the present discussion, in the very low-temperature regime and under the assumption of a large energy gap $\Delta$, the relevant thermodynamic quantities at each stage of the cycle are:
  1. (A)
    Point A ($\lambda_{H},T_{H}$): The internal energy approaches the ground state energy $\epsilon_{0}$, and the entropy corresponds solely to the ground state degeneracy:
    $U_{\lambda_{H}}^{(A)}(T_{H})$ $\simeq\epsilon_{0},$ $S_{\lambda_{H}}^{(A)}(T_{H})=k_{B}\ln\quantity(g_{0}).$ (14)
  2. (B)
    Point B ($\lambda_{\text{crit}},T_{H}$): Despite the higher temperature, the internal energy remains at the critical ground state energy, and the entropy reflects the QCP degeneracy. The thermal energy is insufficient to significantly populate excited states:
    $U_{\lambda_{\text{crit}}}^{(B)}(T_{H})$ $=\epsilon_{\text{ground}}^{\text{(crit)}},$ $S_{\lambda_{\text{crit}}}^{(B)}(T_{H})=k_{B}\ln\quantity(g_{\text{crit}}).$ (15)
  3. (C)
    Point C ($\lambda_{\text{crit}},T_{L}$): The internal energy remains constant at the critical ground state value, and the entropy again reflects the QCP degeneracy:
    $U_{\text{crit}}^{(B)}(T_{H})U_{\lambda_{\text{crit}}}^{(C)}(T_{L})$ $=\epsilon_{\text{ground}}^{\text{(crit)}},$ $S_{\lambda_{\text{crit}}}^{(C)}(T_{L})=k_{B}\ln\quantity(g_{\text{crit}}).$ (16)
  4. (D)
    Point D ($\lambda_{H},T_{L}$): The entropy is determined by the ground state degeneracy at $\lambda_{H}$, and the internal energy approaches $\epsilon_{0}$, as the temperature cannot populate excited states:
    $U_{\lambda_{H}}^{(D)}(T_{L})$ $\simeq\epsilon_{0},$ $S_{\lambda_{H}}^{(D)}(T_{L})=k_{B}\ln\quantity(g_{0}).$ (17)
Thus, using the thermodynamic functions presented in Eqs. (1417), the heat exchanges for each stage of the Stirling cycle are explicitly analyzed:
$Q_{AB}$ $=T_{H}\quantity[S_{\lambda_{\text{crit}}}^{(B)}(T_{H})-S_{\lambda_{H}}^{(A)}(T_{H})]=k_{B}T_{H}[\ln(g_{\text{crit}})-\ln(g_{0})]=k_{B}T_{H}\ln\quantity(\frac{g_{\text{crit}}}{g_{0}}),$ (18)
$Q_{BC}$ $=U_{\lambda_{\text{crit}}}^{(C)}(T_{L})-U_{\lambda_{\text{crit}}}^{(B)}(T_{H})=0,$ (19)
$Q_{CD}$ $=T_{L}\quantity[S_{\lambda_{H}}^{(D)}(T_{L})-S_{\lambda_{\text{crit}}}^{(C)}(T_{L})]=k_{B}T_{L}[\ln(g_{0})-\ln(g_{\text{crit}})]=-k_{B}T_{L}\ln\quantity(\frac{g_{\text{crit}}}{g_{0}}),$ (20)
$Q_{DA}$ $=U_{\lambda_{H}}^{(A)}(T_{H})-U_{\lambda_{H}}^{(D)}(T_{L})=0.$ (21)
Thus, the work performed in the Stirling cycle can be expressed as the sum of the exchanged heats. Assuming that the hot reservoir is set as $T_{H}=T_{L}+\delta$, the net work becomes
$W$ $=k_{B}(T_{H}-T_{L})\ln\quantity(\frac{g_{\text{crit}}}{g_{0}}),$ (22)
$\boxed{W=k_{B}\delta\ln\quantity(\frac{g_{\text{crit}}}{g_{0}})}.$ (23)
where $g_{\text{crit}}$ and $g_{0}$ denote the ground-state degeneracies at the critical and high-field points, respectively, and $\delta$ represents the temperature difference between the reservoirs. By construction, the engine efficiency is identical to the Carnot efficiency.
$\eta_{C}=\frac{\displaystyle{k_{B}T_{H}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}-k_{B}T_{L}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}{\displaystyle k_{B}T_{H}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}=1-\frac{T_{L}}{T_{H}}=\frac{1}{1+\frac{TL}{\delta}}.$ (24)
Notably, while ideal Stirling cycles reach Carnot efficiency only in the presence of a perfect regenerator, in Eq. (24) no regenerator is required. At low temperatures, the thermal population remains confined within the degenerate ground-state manifold, representing a distinct quantum advantage over engines operating with classical working substances.
If the machine operates as a refrigerator, either by reversing the cycle direction or due to intrinsic system properties, the coefficient of performance ($\text{COP}_{R}$) can be defined as an analog to the engine efficiency. However, unlike the efficiency, the COP has no upper bound. For this analysis, the cycle in Fig. 3 is reversed, consequently inverting the labels identifying the stages. The resulting coefficient satisfies the Carnot refrigerator condition:
$COP_{R}$ $=\frac{Q_{\text{out}}}{|W|}=\frac{Q_{DC}+Q_{CB}}{|Q_{AD}+Q_{DC}+Q_{CB}+Q_{BA}|},$ (25)
$COP_{R}$ $=\frac{k_{B}T_{L}\ln\!\left(\tfrac{\displaystyle g_{\text{crit}}}{g_{0}}\right)}{\displaystyle k_{B}T_{H}\ln\!\left(\tfrac{\displaystyle g_{\text{crit}}}{g_{0}}\right)-k_{B}T_{L}\ln\!\left(\tfrac{\displaystyle g_{\text{crit}}}{g_{0}}\right)},$ (26)
$COP_{R}^{(C)}$ $=\frac{T_{L}}{T_{H}-T_{L}}=\frac{T_{L}}{\delta}.$ (27)
Similarly, if the machine operates as a heat pump, the coefficient of performance ($\text{COP}_{HP}$) is defined as the ratio between the heat delivered to the hot reservoir and the net work performed on the system. In our cycle, the heat absorbed by the hot reservoir corresponds to $Q_{AB}$, while the net work is the sum of all heat exchanges. Thus, the ($\text{COP}_{HP}$) can be expressed as:
$\mathrm{COP}_{\!HP}$ $=\frac{Q_{\text{in}}}{|W|}=\frac{Q_{BA}+Q_{AD}}{Q_{DA}+Q_{DC}+Q_{CB}+Q_{BA}},$ (28)
$\mathrm{COP}_{\!HP}$ $=\frac{k_{B}T_{H}\ln\!\left(\tfrac{\displaystyle g_{\text{crit}}}{g_{0}}\right)}{\displaystyle k_{B}(T_{H}-T_{L})\ln\!\left(\tfrac{\displaystyle g_{\text{crit}}}{g_{0}}\right)},$ (29)
$\mathrm{COP}_{\!HP}^{(C)}$ $=\mathrm{COP}_{R}^{(C)}+1=\frac{T_{H}}{T_{H}-T_{L}}=\frac{T_{H}}{\delta}.$ (30)
In the following subsection, the effect of a small energy gap is analyzed, where the first excited state can be thermally occupied even for a small temperature increment $\delta$.
III.3 Thermal Excitations and Small-Gap Effects
To evaluate the effect of higher energy-level occupation on the system’s work and heat exchange, we consider that partial population of the first excited state occurs only when the system is in contact with the hot reservoir at temperature $T_{H}=T_{L}+\delta$ and high fixed parameter $\lambda_{H}$, corresponding to point A in Fig. 3. This contribution is treated as a small perturbation, since the other thermodynamic points remain unaffected. The low temperature $T_{L}$ ensures that only the ground state is populated at point D.
Accordingly, the correction is localized at point A and can be described using an effective two-level model within the canonical ensemble, where the energy gap $\Delta$ in Fig. 1 is sufficiently small to allow partial excitation of the first excited state:
$Z_{\lambda_{H}}(T)$ $=g_{0}\displaystyle e^{-\frac{\epsilon_{0}}{k_{B}T}}+g_{1}e^{-\frac{\epsilon_{0}+\Delta}{k_{B}T}},\qquad\Delta=\epsilon_{1}-\epsilon_{0}.$ (31)
Where $g_{i}$ are the degeneracies of the ground-state $\epsilon_{0}$ an0d the first excited state $\epsilon_{1}$, respectively. Using the standard relations of the canonical ensemble and defining $\Delta$ as the energy gap between the first excited state and the groundstate, the Helmholtz free energy of the system becomes:
$F_{\lambda_{H}}(T)$ $=-k_{B}T\ln\left(Z_{\lambda_{H}}(T)\right)=-k_{B}T\ln\left(e^{-\frac{\displaystyle\epsilon_{0}}{k_{B}T}}\left[g_{0}+g_{1}e^{-\frac{\Delta}{k_{B}T}}\right]\right),$ (32)
$F_{\lambda_{H}}(T)$ $=\epsilon_{0}-k_{B}T\ln g_{0}-\displaystyle k_{B}T\ln\left(1+\frac{g_{1}}{g_{0}}e^{-\frac{\Delta}{k_{B}T}}\right).$ (33)
From the free energy, the internal energy can be calculated, and the entropy can be expressed using the corresponding Legendre transformation for the internal energy as
[…]
Expressing the entropy in Eq. (34) in terms of the internal and Helmholtz free energies is convenient, as only the latter contributes across the cycle. Consequently, the perturbation affects solely the heat exchange at point A. The heat $Q_{AB}$ of the first isothermal process can thus be written in terms of the Helmholtz and internal energies, while the isoparametric process at the high control parameter, associated with the heat $Q_{AD}$, is given by the difference in internal energies at fixed $\lambda_{H}$. At point D, the internal energy corresponds exclusively to the ground state due to the sufficiently low temperature. Thus, the expressions for the heats involved are given by
$Q_{AB}$ $=k_{B}T_{H}\ln\left(g_{\text{crit}}\right)-T_{H}S_{\lambda_{H}}(T_{H}),$ $Q_{AD}=U_{\lambda_{H}}(T_{H})-\epsilon_{0}.$ (35)
Consequently, in the limit $\Delta\gg k_{B}T_{L}$, this term is negligible. Therefore, the heat exchanges for the complete cycle are:
$Q_{AB}$ $=T_{H}\left[S_{\lambda_{\text{cri t}}}^{(B)}(T_{H})-S_{\lambda_{H}}^{(A)}(T_{H})\right]=k_{B}T_{H}\ln\left(g_{\text{crit}}\right)-U_{\lambda_{H}}^{(A)}(T_{H})+F^{(A)}_{\lambda_{H}}(T_{H}),$ (36)
$Q_{BC}$ $=U_{\lambda_{\text{crit}}}^{(C)}(T_{L})-U_{\lambda_{\text{crit}}}^{(B)}(T_{H})=0,$ (37)
$Q_{CD}$ $=T_{L}\left[S_{\lambda_{H}}^{(D)}(T_{L})-S_{\lambda_{\text{crit}}}^{(C)}(T_{L})\right]=k_{B}T_{L}\left[\ln(g_{0})-\ln(g_{\text{crit}})\right]=-k_{B}T_{L}\ln\left(\frac{g_{\text{crit}}}{g_{0}}\right),$ (38)
$Q_{DA}$ $=U_{\lambda_{H}}^{(A)}\left(T_{H}\right)-U_{\lambda_{H}}^{(D)}\left(T_{L}\right)=U_{\lambda_{H}}^{(A)}\left(T_{H}\right)-\epsilon_{0}.$ (39)
The work in the Stirling cycle is calculated as the sum of all heat exchanges. As seen, the internal energy terms cancel, yielding a contribution corresponding to the free energy component:
$W$ $=k_{B}T_{L}\ln\left(\frac{g_{\text{crit}}}{g_{0}}\right)+k_{B}T_{H}\ln\left(g_{\text{crit}}\right)+F_{\lambda_{H}}^{(A)}\left(T_{H}\right)-\epsilon_{0}.$ (40)
By explicitly substituting the free energy from Eq.(33) into Eq.(40), the terms associated with the ground state cancel out, yielding
$W$ $=k_{B}\delta\ln\left(\frac{g_{\text{crit}}}{g_{0}}\right)-k_{B}(T_{L}+\delta)\ln\left(1+\frac{g_{1}}{g_{0}}\displaystyle e^{-\frac{\Delta}{k_{B}(T_{L}+\delta)}}\right).$ (41)
It is observed that, in the limit of a large energy gap, the previous result from Eq. (24) is recovered with an additional contribution at the Carnot limit. However, the thermal population of the first excited state reduces the useful work. To investigate its impact, the system efficiency is now analyzed from Eq. (13) substituting the heat terms from Eqs. (3639). The resulting expression can be written as follows
$\eta=1-\frac{\displaystyle k_{B}T_{L}\ln\left(\frac{\displaystyle g_{\text{crit}}}{g_{0}}\right)}{\displaystyle k_{B}T_{H}\ln\left(\frac{g_{\text{crit}}}{g_{0}}\right)-k_{B}T_{H}\ln\left(1+\frac{g_{1}}{g_{0}}e^{-\frac{\Delta}{k_{B}T_{H}}}\right)}.$ (42)
To account for the correction introduced by the heat term $Q_{AD}$ to Carnot efficiency, Eq. (24) is expanded as a power series. Since the critical point involves the degeneracies $g_{0}$ and $g_{1}$, the condition $\ln\!\left(1+\frac{g_{1}}{g_{0}}e^{-\Delta/(k_{B}T_{H})}\right)<\ln\!\left(\frac{g_{\text{crit}}}{g_{0}}\right)$ is always satisfied, enabling a geometric expansion. The result can then be expressed in terms of the Carnot efficiency as
$\eta$ $=\eta_{C}-\frac{T_{L}}{T_{H}}\sum_{n=1}^{\infty}\quantity(\frac{\displaystyle\ln\left(1+\frac{g_{1}}{g_{0}}e^{-\frac{\Delta}{k_{B}T_{H}}}\right)}{\displaystyle\ln\left(\frac{g_{\text{crit}}}{g_{0}}\right)})^{n}.$ (43)
Inclusion of higher thermal populations reduces the maximum attainable efficiency, keeping it strictly below the Carnot limit. This reduction arises because part of the thermal energy is expended in populating excited states rather than contributing to useful work. In the limit of a large energy gap and/or sufficiently low temperature, these corrections vanish, and the Carnot efficiency is asymptotically recovered.
IV General Framework for Quantum Stirling Cycles: Validation and Application to Critical Systems with Fibonacci-Type Degeneracies
Having established the theoretical framework for both ideal critical cycles and systems with a small energy gap, we now validate these models against quantum systems with QCPs reported in the literature. This includes quantum critical systems, quantum heat engines that reach Carnot efficiency, and exactly solvable models. We examine a range of systems with diverse interactions to validate the derived formulas through postdictions of previously reported results. Furthermore, we provide new predictions for systems exhibiting nontrivial degeneracies, such as the critical point degeneracies in open and closed quantum Ising chains, which are characterized by Fibonacci and Lucas numbers.
IV.1 Validation Across Reported Quantum Systems
IV.1.1 Two spin-1/2 systems
Two-qubit systems have been widely studied in condensed matter and quantum information literature, often with different types of interactions. To provide a unified framework for these cases, we introduce a generalized XXZ Hamiltonian that collects the most relevant contributions, written in natural units of the exchange constant $J$:
$\hat{\mathcal{H}}_{2\,spin-1/2}$ $=J(\hat{S}_{1}^{x}\hat{S}_{2}^{x}+\hat{S}_{1}^{y}\hat{S}_{2}^{y})+J_{z}\hat{S}_{1}^{z}\hat{S}_{2}^{z}+B_{z}\hat{S}_{\text{tot }}^{z}+K_{y}(\hat{S}_{\text{tot }}^{y})^{2}+D\left(\hat{\mathbf{S}}_{1}\times\hat{\mathbf{S}}_{2}\right)\cdot\hat{\mathbf{z}}$. (44)
The Hamiltonian in Eq. (44) incorporates several fundamental interactions within the standard spin-operator formalism. The antiferromagnetic exchange constant $J$ couples nearest-neighbor spins through a Heisenberg XX interaction, where the spin operators $\hat{\bm{S}}_{i}=(\hat{S}_{i}^{x},\hat{S}_{i}^{y},\hat{S}_{i}^{z})$ satisfy the usual angular momentum commutation relations. The total spin components are defined as $\hat{S}_{\text{tot}}^{\alpha}=\sum_{i=1}^{N}\hat{S}_{i}^{\alpha}$ with $\alpha\in\{x,y,z\}$. The $J_{z}$ term couples the $z$-components of both spins, and for $J_{z}=J$ the system reduces to the isotropic Heisenberg XXX model. The Zeeman term, proportional to $B_{z}$, corresponds to a homogeneous magnetic field applied along the $z$ axis. The uniaxial magnetic anisotropy $K_{y}$ couples to the square of the total $y$-component of spin. Finally, the Dzyaloshinskii–Moriya interaction (DMI), parameterized by $D$, introduces an antisymmetric exchange that breaks inversion symmetry.
By selectively enabling or suppressing interaction terms, previously reported two-qubit models can be recovered as limiting cases of the general Hamiltonian. This construction provides a unified framework within which our formula can be benchmarked against diverse quantum configurations operating as Stirling engines. The corresponding parameter sets and the quantitative agreement are summarized in Table 2.
Interaction
Mobile
param.
$\lambda_{L}$ $\lambda_{H}$ $T_{L}$ $T_{H}$ $g_{\text{crit}}$ $g_{0}$ $W_{\text{sim}}$ $W_{\text{formula}}$ Error
J=1&D=0J=1\;\&\;D=0
$K_{y}=J_{z}=0$
Purkait et al. [29]
$B_{z}$ 1 3 0.15 0.25 2 1 0.0693147 $0.1\ln 2$ $4.0\times 10^{-7}$
$J=1,\;K_{y}=0.5$
$J_{z}=D=0$
Araya et al. [33]
$B_{z}$ $\sqrt{2}$ 3 0.12 0.20 2 1 0.0693147 $0.08\ln 2$ $1.6\times 10^{-5}$
$J=1,\;K_{y}=0$
$2D=J_{z}=1$
Pili et al. [35]
$B_{z}$ 1.06 4 0.10 0.20 2 1 0.0693161 $0.1\ln 2$ $2.0\times 10^{-5}$
$B=D=0$
$K_{y}=J_{z}=0$
Cruz et al. [26]
$J$ -2.75 -1 20 25 4 1 6.90505 $5\ln 4$ $7.7\times 10^{-3}$
Table 2: Comparison of the generalized two-qubit Hamiltonian under different interaction configurations, operating as a quantum Stirling engine. The table lists the parameter sets considered, together with the corresponding degeneracies, simulated work output ($W_{\text{sim}}$), and the predictions of the analytical expression ($W_{\text{formula}}$). Temperatures and work are given in units of $k_{B}$, while the mobile parameter is expressed in exchange-constant units.
Table 2 summarizes several interaction configurations of the Hamiltonian in Eq.(44) along with their corresponding references. The mobile parameter varies between models, though the magnetic field is most commonly employed as the control variable. Thermal reservoirs are set at low and closely spaced temperatures. The simulated work output $W_{\text{sim}}$, given in units of $k_{B}$, is obtained from full numerical calculations and compared with the analytical prediction of Eq.(23), which depends solely on degeneracies and temperature differences for systems operating at Carnot efficiency.
In the case of the pure Heisenberg XX model, Purkait and Biswas [29] identified a QCP at $B=1$ with a twofold critical degeneracy. For $B=3$, where the ground state is non-degenerate, and temperatures $T_{L}=0.15$ and $T_{H}=0.25$, the system reaches Carnot efficiency, and the simulated work matches precisely the prediction of Eq. (23).
Introducing a planar anisotropy of $K_{y}=0.5$, Araya et al. [33] showed that the critical field shifts to $B\simeq\sqrt{2}$ while preserving the twofold degeneracy. With $B=3$ as the high-field parameter and temperatures $T_{L}=0.12$ and $T_{H}=0.20$, the work remains proportional to $\ln 2$, again in perfect agreement with the formula.
The combined effect of a homogeneous field and a Dzyaloshinskii–Moriya interaction was analyzed by Pili et al. [35]. For $D=1/2$, the critical point shifts to $B=1.06$, while the ground-state degeneracy remains unchanged. Operating the cycle with $B=4$, $T_{L}=0.10$, and $T_{H}=0.20$, the work continues to follow Eq. (23), demonstrating its robustness even away from the strict critical regime.
Cruz et al. [26] investigated an experimental dinuclear metal complex modeled as a two-qubit system with only nearest-neighbor interactions. Measurements of work, absorbed and released heat, and efficiency were found to be consistent with theoretical predictions. For the purposes of this manuscript, parameters around $J_{L}\simeq-2.75$ with reservoir temperatures $T_{L}=20$ K and $T_{H}=25$ K are considered. To further approach the critical regime at $J=0$, the interaction is extended down to $J_{H}=-1$, slightly beyond the values reported in the experiment. In this range, the work output approaches the Carnot limit, though the deviation from Eq. (23) remains on the order of $10^{-3}$ due to the high-field point not coinciding exactly with the fourfold critical degeneracy. Extending the parameter closer to $J\simeq 0$ restores agreement with errors as small as $10^{-10}$, although such conditions would not correspond to physically realizable parameters in an endoreversible cycle.
IV.1.2 Nth spin-s system
The two–qubit benchmarks establish quantitative agreement between the analytical expression in Eq. (23) and simulated work across distinct interaction channels. The next step is to probe robustness beyond the minimal local Hilbert space and to assess the role of topology. This requires moving to higher on–site spin and to extended geometries (open chains vs. rings), for which a compact many–body description is provided below.
Recent studies demonstrate that topology, system size, and on-site spin magnitude significantly influence quantum thermal machine performance. To capture these effects systematically, we introduce a generalized Heisenberg-type Hamiltonian for chains of NN spin-$s$ sites:
$\hat{\mathcal{H}}_{\mathrm{spin-s}}^{(N)}$ $=J\sum_{i=1}^{N-1}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{i+1}+B\hat{S}_{\mathrm{tot}}^{z}+K_{y}\left(\hat{S}_{\mathrm{tot}}^{y}\right)^{2}+K_{z}\sum_{i=1}^{N}\left(\hat{S}_{i}^{z}\right)^{2}.$ (45)
This Hamiltonian (45) integrates the key interactions previously established: Heisenberg XXX exchange coupling ($J$), Zeeman field ($B$), easy-plane anisotropy ($K_{y}$), and single-ion anisotropy ($K_{z}$) that promotes alignment of individual spins along the $z$-axis. When the summation is extended to $N$ terms and periodic boundary conditions are imposed ($\hat{\mathbf{S}}_{N+1}\equiv\hat{\mathbf{S}}_{1}$), the Hamiltonian can describe either open chains or closed rings, enabling systematic exploration of topological effects on work extraction and efficiency in quantum Stirling cycles.
This framework provides a unified foundation for analyzing how spin magnitude ($s$), geometry, and interaction symmetry shape the performance of quantum thermal devices. Table 3 compares representative cases of quantum Stirling engines with the analytical expression in Eq. (23) at Carnot efficiency. The results show consistent agreement between simulations and theory across different geometries, interaction types, and system sizes, with the magnetic field acting as the control parameter and thermal reservoirs maintained at low, closely spaced temperatures.
Interaction
Mobile
param.
$\lambda_{L}$ $\lambda_{H}$ $T_{L}$ $T_{H}$ $g_{\text{crit}}$ $g_{0}$ $W_{\text{sim}}$ $W_{\text{formula}}$ Error
$J=1\;\&\;\textbf{Chain}$
$K_{y}=-0.44$
$K_{z}=0$
Castorene et al. [34]
(N = 3, Spin-1/2)
$B$ 2.412 3.5 0.25 0.15 2 1 0.0692511 $0.1\ln 2$ $9.1\times 10^{-4}$
$J=1\;\&\;\textbf{Ring}$
$K_{y}=-0.44$
$K_{z}=0$
Castorene et al. [34]
(N = 3, Spin-1/2)
$B$ 2.412 3.5 0.25 0.15 3 1 0.109734 $0.1\ln 3$ $1.1\times 10^{-3}$
$J=1\;\&\;\textbf{Chain}$
$\mathbf{K_{z}=-2/3}$
$K_{y}=0$
Castorene et al. [24]
(N = 2, Spin-1)
$B$ 4/3 3.0 0.10 0.05 3 1 0.054973 $0.05\ln 3$ $7.7\times 10^{-4}$
$J=1\;\&\;\textbf{Chain}$
$\mathbf{K_{z}=-1/10}$
$K_{y}=0$
Castorene et al. [24]
(N = 2, Spin-1)
$B$ 1.90 3.0 0.10 0.05 2 1 0.0346646 $0.05\ln 2$ $2.0\times 10^{-4}$
Table 3: Comparison of the generalized spin-$s$ chain Hamiltonian (45) for different interaction configurations and system sizes $N$, operating as quantum Stirling engines. The table reports the parameter sets, critical degeneracies, simulated work output ($W_{\text{sim}}$), and the predictions of Eq. (23), showing quantitative agreement with results reported in the literature. Temperatures and work are given in units of $k_{B}$, while the mobile parameter is expressed in exchange-constant units.
Castorene et al. [34] investigated isotropic Heisenberg XXX chains and rings of three qubits with a uniaxial anisotropy along the $y$-axis, using the magnetic field as the control parameter. As discussed earlier, such anisotropy shifts the critical field without altering the degeneracy structure. For a critical field at $B\simeq 2.41$ and a high-field point at $B=3.5$, with reservoirs at $T_{L}=0.15$ and $T_{H}=0.25$, the work output followed Eq. (23) in both geometries. Notably, the ring configuration exhibited a larger work output, which can be attributed to the additional translational symmetry at zero field, yielding one extra critical degeneracy compared to the open chain.
In a different study, Castorene et al. [24] analyzed isotropic spin-1 dimers with single-ion anisotropy along the $z$-axis and an external magnetic field of opposite sign, which introduces frustration in the energy levels. Specific anisotropy values induce either a triple degeneracy point (TDP) at $K_{z}=-2/3$ or double degeneracy points (DDPs) at $K_{z}=-1/10$, each associated with critical magnetic fields. For instance, at $B=4/3$ the system exhibits a TDP, while at $B\simeq 1.90$ one of the DDPs emerges. Stirling cycles performed with $B_{H}=3.0$ and reservoirs at $T_{L}=0.05$ and $T_{H}=0.10$ yielded work outputs proportional to $\ln 3$ and $\ln 2$, respectively, in full agreement with Eq. (23).
Overall, these results highlight the robustness of the proposed framework across diverse quantum critical scenarios. The analytical expression for the work remains valid not only for two-qubit systems with additional interactions such as Dzyaloshinskii–Moriya coupling or planar anisotropy, but also for extended configurations involving higher spins, distinct topologies (chains and rings), and complex local anisotropies. This universality underscores the broad applicability of Eq. (23) to quantum thermal devices operating near criticality.
To further demonstrate the generality of the approach, an additional analysis of a non-Heisenberg spin model, the Lipkin–Meshkov–Glick (LMG) model, was carried out in Appendix A. The results confirm that the proposed formula remains accurate even for systems characterized by collective spin interactions and more intricate energy landscapes.
IV.2 Quantum Ising model at criticality with Fibonacci-Lucas Degeneracies
The scalability of the work formula can be addressed by considering systems with an increasing number of interacting particles. A natural case study is the quantum antiferromagnetic Ising model in the $z$ direction, one of the simplest many-body systems that exhibits a non-trivial QCP. The model is studied under a longitudinal magnetic field applied along the $z$ axis, which drives the transition and enables exploration of how critical degeneracies scale with system size. The Hamiltonian of this system is given by
$\hat{\mathbf{\mathcal{H}}}$ $=J\sum_{i=1}^{N-1}{S}^{z}_{i}{S}_{i+1}^{z}+B\sum_{i=1}^{N}{S}^{z}_{i}.$ (46)
This Hamiltonian describes a chain of $N$ qubits with nearest–neighbor antiferromagnetic interactions. The exchange constant $J$ denotes the isotropic antiferromagnetic coupling constant between adjacent spins $(J>0)$, while $\hat{S}^{z}_{i}$ represents the Pauli operator along the $z$ direction. The longitudinal external magnetic field $B$ is applied along the positive $z$ axis and, in physical units, is given by $B=\mu_{B}g_{z}H$, where $\mu_{B}$ is the Bohr magneton and $g_{z}$ the Landé $g$-factor. For convenience, $B$ is expressed in units of the exchange constant. Extending the summation to $N$ terms with the identification $\hat{S}^{z}_{N+1}=\hat{S}^{z}_{1}$ imposes periodic boundary conditions, thereby mapping the open chain system onto a closed spin ring.
In this framework, the system exhibits a QCP at finite field, which occurs when the ratio $B/J=1$. Such criticality arises for both open chains and closed rings of arbitrary length $N\geq 3$.
Correspondingly, the ground-state energy at the critical field and beyond, with respect to the magnetic field, is given by [40]
$\epsilon_{\text{Ising}}^{(\text{ground})}=\frac{J}{4}(N-1)-\frac{NB}{2},\qquad\frac{J}{B}\geq 1.$ (47)
For periodic boundary conditions, the factor accompanying the exchange constant $J$ extends as $(N-1)\rightarrow N$, while the magnetic field term remains unchanged. This implies that the open-chain geometry possesses a lower ground-state energy than the closed ring.
Accordingly, the Stirling cycle shown in Fig. 2 is implemented by driving the system from the quantum critical point at $\lambda_{L}=B/J=1$ to a higher magnetic field $\lambda_{H}=B/J=3$. The process is simulated within the canonical ensemble over the complete set of $2^{N}$ energy levels for both the open-chain and closed-ring geometries. This framework allows a direct comparison between the numerically computed work outputs $(W_{\text{Ising}}^{\text{Chain}},W_{\text{Ising}}^{\text{Ring}})$ and the analytical prediction from Eq. (23), valid for arbitrary system size $N$. To establish this connection, it is necessary to determine how the critical ground-state degeneracy associated with the ground-state energy of Eq. (47) scales with $N$. In this context, recent studies [40, 41] showed that, for one-dimensional antiferromagnetic Ising model, the degeneracy at the QCP follows the Fibonacci sequence $F_{N}$, defined recursively as $F_{N}=F_{N-1}+F_{N-2}$ with $F_{0}=F_{1}=1$. In contrast, periodic boundary conditions give rise to a nontrivial degeneracy governed by the Lucas sequence, which can be expressed in terms of the Fibonacci numbers as $L_{N}=F_{N+1}+F_{N-1}$. Using Binet’s formula, both sequences can be written in closed form through the golden ratio $\varphi=(1+\sqrt{5})/2$:
$F_{N}=\frac{\varphi^{N}-(-\varphi^{-1})^{N}}{\sqrt{5}},\qquad L_{N}=\varphi^{N}+(-1)^{N}\varphi^{-N}.$ (48)
Figure 4: Quantum Stirling work $W$ [meV] as a function of the number of sites NN for the antiferromagnetic Ising chain and ring ($J=1$), evaluated at the QCP $\lambda_{L}=B/J=1$ and a high-field parameter $\lambda_{H}=B/J=3$, with reservoir temperatures $T_{L}=0.1$ and $T_{H}=0.2$. In all cases the system operates at Carnot efficiency $\eta_{C}=0.5$. Red (black) dots represent simulated cycles for the chain (ring) system. Yellow and green markers show the predictions of Eqs. (49-50) based on Fibonacci and Lucas degeneracies at the QCP, respectively, while blue (orange) lines correspond to the large-NN asymptotic expressions in Eqs. (51-52).
Thus, performing the Stirling cycle from the QCP at $B/J=1$, where the Fibonacci–Lucas degeneracies of Eq. (48) apply, to a high-field configuration at $B/J=3$ with a non-degenerate ground state yields
$W_{\text{Formula}}^{\text{Chain}}$ $=k_{B}\delta\ln\quantity(\frac{\varphi^{N}-\left(-\varphi^{-1}\right)^{N}}{\sqrt{5}}),$ (49)
$W_{\text{Formula}}^{\text{Ring}}$ $=k_{B}\delta\ln\quantity(\varphi^{N}-\left(-\varphi^{-1}\right)^{N}).$ (50)
These previous expressions can be treated for the case of small system sizes $N$, observing that the work of both the chain and the ring does not scale linearly with $N$, reflecting a genuinely quantum effect that contrasts with the classical thermodynamic expectation of extensivity. Specifically, Eq. (49) shows that the work for the chain is proportional to the $N$-th Fibonacci number, while for Eq. (50) depict that the ring follows the $N$-th Lucas number, with the chain producing less work.
In the asymptotic limit $N\gg 1$, the inverse golden-ratio $(\varphi^{-N})$ terms can be neglected, yielding the approximate expressions
$W_{\text{Formula}}^{\text{Chain}}$ $\xrightarrow[N\gg 1]{}Nk_{B}\delta\ln\quantity(\varphi)-k_{B}\delta\ln\sqrt{5},$ (51)
$W_{\text{Formula}}^{\text{Ring}}$ $\xrightarrow[N\gg 1]{}Nk_{B}\delta\ln\quantity(\varphi).$ (52)
Eqs. (51-52) show that both cases scale linearly with $N$, but the chain includes a constant offset of $\ln\sqrt{5}$, highlighting a residual distinction between quantum thermodynamics and its classical counterpart. However, in the thermodynamic limit $N\ggg 1$, the constant term becomes negligible, and both chain and ring topologies recover classical extensivity, yielding the same work output determined solely by the system size. This behavior is consistent with results reported in the literature for the classical Ising model.
Finally, Fig. 4 summarizes the main results, displaying the quantum Stirling work W as a function of the system size N for both open-chain and closed-ring configurations of the antiferromagnetic Ising model. The numerical simulations (red and black markers) exhibit excellent agreement with the analytical predictions derived from the Fibonacci and Lucas degeneracies (yellow and green markers), as well as with the large-N asymptotic limits (blue and orange lines). This consistency across sizes and topologies demonstrates the robustness and general validity of the proposed theoretical framework.
V Conclusions
We have performed a comprehensive thermodynamic analysis of quantum systems exhibiting a QCP induced by a control parameter at zero temperature, using the canonical ensemble formalism. From these quantities, a quantum Stirling cycle was constructed under sufficiently low temperatures such that only the ground state remains significantly populated. The cycle begins at the QCP, characterized by critical degeneracy, and ends at a higher control parameter where the ground state becomes energetically isolated from the excited manifold.
This framework yields analytical expressions for the finite net work and Carnot-like efficiency in terms of the temperature difference between the thermal reservoirs and the ratio between the degeneracies of the critical and isolated ground state. The role of finite-temperature effects was also examined, revealing that partial thermal occupation of excited states reduces both useful work and maximum efficiency due to energy redistribution within the system.
The proposed work expression was validated through postdictions and predictions across diverse spin-$s$ Heisenberg systems, incorporating Dzyaloshinskii–Moriya interactions, uniform magnetic fields, uniaxial anisotropies, and single-ion anisotropies. These results confirm the robustness of the formulation for systems operating near the Carnot regime.
Finally, the extensivity of the work output was analyzed for one-dimensional antiferromagnetic Ising systems with open-chain and closed-ring topologies of length $N$. The work was found to scale with the logarithm of the $N$th Fibonacci number for open chains and with the logarithm of the $N$th Lucas number for rings, indicating a higher useful work in closed geometries for small $N$. In the thermodynamic limit, both configurations converge to a linear dependence on $N$, thereby recovering the classical extensivity of the work.
Acknowledgements.
Authors acknowledge partial support from ANID CIA 250002, Fondecyt Grant (Chile) under contracts 1230055, 1240582 and 1250173. B.C. acknowledges support from Pontificia Universidad Católica de Valparaíso (PUCV) and the Dirección de Postgrado of Universidad Técnica Federico Santa María (UTFSM). B.C. also acknowledges the support of ANID Becas/Doctorado Nacional 21250015.
Author Contributions
Bastian Castorene  — Conceptualization (lead), Data curation (lead), Formal analysis (lead), Investigation (lead), Methodology (lead), Visualization (lead), Writing – original draft (lead), Writing – review & editing (lead).
Martin HvE Groves  — Formal analysis (equal), Investigation (supporting), Validation (equal), Writing – review & editing (lead).
Francisco J. Peña  — Investigation (supporting), Supervision (equal), Validation (lead), Writing – review & editing (lead).
Eugenio E. Vogel  — Conceptualization (supporting), Funding acquisition (lead), Supervision (equal), Validation (equal), Writing – review & editing (supporting).
Patricio Vargas  — Conceptualization (lead), Methodology (lead), Supervision (lead), Writing – original draft (supporting), Writing – review & editing (equal).
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Appendix A Applications in the Lipkin–Meshkov–Glick model
The Lipkin–Meshkov–Glick (LMG) model constitutes a paradigmatic framework in quantum many-body physics, extensively investigated in various experimental platforms such as Bose–Einstein condensates and trapped-ion systems [25]. Recent studies have shown that a quantum Stirling cycle employing the LMG model as its working substance can, under appropriate conditions, reach Carnot efficiency.
The model describes a system of $N$ collectively interacting spins subjected to a uniform magnetic field, incorporating both symmetric and anisotropic spin–spin couplings. Its general Hamiltonian can be written as
$\hat{\mathcal{H}}_{\text{LMG}}$ $=\varepsilon\,S_{z}+V\left(S_{x}^{2}-S_{y}^{2}\right)+W\left(S^{2}-S_{z}^{2}\right),$ (53)
where $S_{\alpha}=\tfrac{1}{2}\sum_{i=1}^{N}\sigma_{i}^{\alpha}$ denotes the collective spin operators, and $\sigma_{i}^{\alpha}$ are the Pauli matrices acting on the $i$-th spin. The first term represents the Zeeman interaction with an external magnetic field of strength $\varepsilon$, while $S^{2}=S_{x}^{2}+S_{y}^{2}+S_{z}^{2}$ corresponds to the total spin magnitude. The coefficients $V$ and $W$ quantify the anisotropic components of the collective spin–spin coupling.
It is convenient to define the following dimensionless control parameters:
$\lambda=-\frac{\varepsilon}{2},\qquad\frac{\lambda_{0}}{N}=-\frac{W+V}{2},\qquad\gamma=\frac{W-V}{W+V}.$ (54)
In the isotropic case $\gamma=1$, the Hamiltonian simplifies to
$\hat{\mathcal{H}}_{\text{LMG}}=-\frac{2\lambda_{0}}{N}\left(S^{2}-S_{z}^{2}\right)-2\lambda S_{z}.$ (55)
By setting $\lambda_{0}=1$, the energy eigenvalues take the compact form
$E(M)=\frac{2}{N}\left(M-\frac{N\lambda}{2}\right)^{2}-\frac{N}{2}\left(1+\lambda^{2}\right)-1,$ (56)
where $M$ assumes discrete values $M\in[-N/2,\,N/2]$.
This collective-spin model fulfills the general requirements for a quantum working substance discussed in the main text. Accordingly, the thermodynamic formalism developed earlier can be directly applied. For $N=3$ and $N=4$ the corresponding quantum Stirling cycles are constructed from the energy spectrum given in Eq. (56), where the parameter $\lambda$ acts as the external thermodynamic variable controlling the energy-level structure throughout the cycle.
Mobile param. $\lambda_{L}$ $\lambda_{H}$ $T_{L}$ $T_{H}$ $g_{\text{crit}}$ $g_{0}$ $W_{\text{sim}}$ $W_{\text{formula}}$ Error
$\lambda$
(N = 3)
0.66 2.5 0.10 0.20 2 1 0.0694418 $0.1\ln 2$ $1.83\times 10^{-3}$
$\lambda$
(N = 4)
0.75 4.0 0.15 0.30 2 1 0.0940877 $0.15\ln 2$ $4.77\times 10^{-2}$
Table 4: Performance of the Lipkin–Meshkov–Glick (LMG) model operating as a quantum Stirling engine for different parameter sets and system sizes $N=3$ and $N=4$, respectively. The table lists the control parameters $\lambda_{L}$ and $\lambda_{H}$, the reservoir temperatures $T_{L}$ and $T_{H}$, the critical degeneracy $g_{\text{crit}}$, the ground-state degeneracy $g_{0}$, and the corresponding work output obtained from simulations ($W_{\text{sim}}$) compared with the analytical prediction of Eq. (23). Quantities are expressed in units of $k_{B}$ for temperature and work, while $\lambda$ is given in dimensionless form. Results show quantitative agreement with Ref. [25].
The excellent agreement between the simulated and analytical results, evidenced by the negligible relative error, confirms the consistency and reliability of our approach. Moreover, for the parameter sets presented in Table 4, the quantum Stirling engine based on the Lipkin–Meshkov–Glick model operates at the Carnot efficiency limit, demonstrating the thermodynamic optimality of the system under the specified conditions. This result further highlights the versatility of the proposed formulation, which remains valid beyond Heisenberg-type spin models and successfully extends to collectively interacting systems such as the LMG model.
Acknowledgements.
Authors acknowledge partial support from ANID CIA 250002, Fondecyt Grant (Chile) under contracts 1230055, 1240582 and 1250173. B.C. acknowledges support from Pontificia Universidad Católica de Valparaíso (PUCV) and the Dirección de Postgrado of Universidad Técnica Federico Santa María (UTFSM). B.C. also acknowledges the support of ANID Becas/Doctorado Nacional 21250015.